# BayesSDF: Surface-Based Laplacian Uncertainty Estimation for 3D Geometry with Neural Signed Distance Fields

### v2

@@ 5.2 Distortion Fields
The distortion field $f_{\mathrm{hash}}(\mathbf{x})$ captures spatial distortions in the SDF field, serving as a basis for uncertainty estimation. To represent the deformation parameters $\mathbf{d}(\mathbf{x})$ on a continuous domain, we employ a HashEncoding that functions as a voxel-like grid implemented through a hash table $\mathbf{d}(\mathbf{x})=f_{\mathrm{hash}}(\mathbf{x})\in\mathbb{R}^{3}$ where $f_{\mathrm{hash}}$ is a function learned by a single-level hash encoding with: $\mathrm{min\_res}=\mathrm{max\_res}=2^{\ell}$ for level-of-detail $\ell$, a hash table size of $2^{3\ell+1}$ entries, linear interpolation between grid corners, and 3 output channels per queried grid vertex. Effectively, for a given point $\mathbf{x}$, the hash encoding looks up an 8-corner neighborhood around the voxel containing $\mathbf{x}$, interpolates among these corners (each storing a 3D deformation vector), and outputs a single 3D vector $\mathbf{d}(\mathbf{x})$. This approach is parameter-efficient while remaining flexible. We denote $\mathbf{aabb}\in\mathbb{R}^{6}$ as the scene bounding box, which is used to normalize $\mathbf{x}$ to a unit cube or appropriate domain before querying the hash grid. Specifically, we generate a ray bundle from the dataset’s rays, containing origins, directions, and near/far bounds. We then use Neus’ hierarchical or “proposal” sampler to refine sample locations along each ray, focusing more samples near scene geometry (i.e., near the SDF zero level set). Let $\{\mathbf{x}_{i}\}$ be the set of sampled 3D points on a ray. When each ray sample $\mathbf{x}\in\mathbb{R}^{3}$ is generated, we transform $\mathbf{x}$ into normalized coordinates $\tilde{\mathbf{x}}=\mathrm{normalize}(\mathbf{x},\mathbf{aabb})$. Then, $\mathbf{d}(\mathbf{x})\;=\;\sum_{c=1}^{8}\alpha_{c}(\tilde{\mathbf{x}})\,\mathbf{d}_{c}\quad\text{where}\quad\bigl{\{}\mathbf{d}c\bigr{\}}{c=1}^{8}\;$ are are the corner deformation vectors that are found from a grid cell index in each dimension (one for each binary combination of $\lfloor\cdot\rfloor or\lceil\cdot\rceil$ in (x,y,z)) and $\alpha_{c}(\tilde{\mathbf{x}})$ are the trilinear interpolation coefficients $(\alpha_{c}(\mathbf{x}i)\;\in\;[0,1],\quad c=1,\ldots,8,)$ such that $\sum{c=1}^{8}\alpha_{c}(\mathbf{x}_{i})=1$. These coefficients are used to blend the deformation vectors from the 8 corners. This procedure is vital not only for generating $\mathbf{d}(\mathbf{x}_{i})$, but also for accumulating partial derivatives. Because each point’s deformation influences eight corner entries, we track which corners are involved and in what proportion.
@@ 5.3 Calculating the Hessian
$\Delta(\mathbf{x})\;=\;\sum_{d=1}^{3}\Bigl{(}\frac{\partial r}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial g}{\partial\mathbf{d}_{d}}^{2}\textbullet\frac{\partial b}{\partial\mathbf{d}_{d}}^{2}\Bigr{)},$
@@ 5.5 Uncertainty Approximation
We interpret these aggregated squared gradients in ‘hessian‘ as proxies for uncertainty. High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color, suggesting the reconstruction at that location is sensitive—and potentially uncertain. Conversely, near-zero partial derivatives indicate that changes in local deformation parameters barely affect the color, suggesting more stable or confident estimates. Given the final accumulated $\mathbf{H}$ stored in our 3D grid, we define an uncertainty function $\sigma$ over the domain:
@@ 6.2 Evaluation Metrics
Figure 3: Africa Dataset: Change in mean absolute error ($\Delta$MAE) as a function of sparsification.
Figure 4: Basket Dataset: Sparsification curve showing the change in mean absolute error ($\Delta$MAE) with increasing levels of sparsification.
@@
    where $\hat{d}(u,v)$ is the predicted depth, $d^{(}u,v)$ is ground-truth depth, and $\Omega(p)$ is the set of remaining pixels after removing a fraction p of the domain based on uncertainty sorting. Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors. If the model correctly identifies ”most uncertain” pixels as having high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
@@
The figures show three colored lines for each dataset/view: the purple line represents sorting by BayesSDF’s uncertainty in descending order, the blue line represents sorting by variance or alternate uncertainty measure (e.g., an ensemble-based variance), and the red line represents the difference between the two curves’ $\Delta\text{MAE}$.
When the red line reaches a low value, it implies that our Hessian-based uncertainty is identifying the same high-error pixels that the variance-based approach finds.
@@ 6.3 Analysis
[…]
@@
  • Despite not always having the absolute lowest metric, BayesSDF’s performance remains highly competitive while avoiding the large computational expense of ensemble training.
@@
    Compact Integration into NeuS: BayesSDF only requires storing a Hessian-based statistic in the resolution-limited deformation grid, leading to a small parameter footprint while seamlessly integrating into the existing NeuS rendering pipeline.
To validate that BayesSDF’s uncertainty maps correspond to poor geometry, we compare high-uncertainty regions against ground-truth geometry across all datasets. We observe that high Hessian-based uncertainty consistently aligns with areas exhibiting large depth errors or missing surface detail such as reflective regions, occlusions, or sparse view coverage. These findings suggest that our Laplace approximation effectively identifies instability in the underlying SDF geometry, making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
@@ 7 Applications
The BayesSDF framework offers a transformative approach to uncertainty quantification, with significant implications for several domains that require precise and reliable 3D reconstructions. Below, we discuss the core applications of BayesSDF, emphasizing its technical strengths and relevance across various fields. BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics. BayesSDF quantifies spatial ambiguities in reconstructed regions. These uncertainty maps enable robotic systems to detect areas of high ambiguity, invaluable for autonomous navigation in dynamic or cluttered environments. For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements. Physics-based simulations in fields such as engineering and medical imaging rely on accurate 3D models to simulate phenomena such as fluid dynamics or stress distribution. BayesSDF enhances these simulations by providing geometric models with quantified uncertainty. For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools. BayesSDF opens new avenues in scientific research fields that depend on precise 3D modeling, such as computational biology and geology. For example, in paleontology, BayesSDF can assist in reconstructing fossilized remains with quantified confidence levels, allowing researchers to distinguish between highly reliable and uncertain reconstructions.
@@ 8 Conclusion
We introduce BayesSDF, a probabilistic framework designed for uncertainty quantification in SDF-based 3D reconstruction models. Motivated by scientific and physical simulation applications where accurate surface geometry is paramount, BayesSDF integrates a Laplace approximation technique over signed distance fields to quantify local reconstruction instability. Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs, and our method builds on this advantage to provide meaningful geometric uncertainty. The result is a compact, efficient uncertainty estimator that aligns uncertainty directly with surface quality. By utilizing models like NeuS’ volumetric SDF space, BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification. The technical contributions of BayesSDF include a deformation field model based on hierarchical hashing, a computationally efficient Hessian-based curvature estimation, and ensemble-based epistemic uncertainty metrics. Together, these components address critical challenges in 3D uncertainty estimation, including the misalignment of uncertainty with surface geometry and the high computational costs associated with traditional sampling-based approaches. Through rigorous experiments on synthetic and real-world datasets, BayesSDF demonstrates outstanding performance, achieving very high calibration results and reduced computational overhead. Looking forward, the principles of BayesSDF can be extended to broader applications, such as dynamic scene understanding and hybrid models that combine SDFs with density-based neural representations. Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks presents exciting opportunities for real-time 3D uncertainty quantification.

### v3

@@ 4.2 Distortion Fields
[…]
@@ 4.3 Calculating the Hessian
$\Delta(\mathbf{x})\;=\;\sum_{d=1}^{3}\left(\left(\frac{\partial r}{\partial\mathbf{d}_{d}}\right)^{2}\cdot\left(\frac{\partial g}{\partial\mathbf{d}_{d}}\right)^{2}\cdot\left(\frac{\partial b}{\partial\mathbf{d}_{d}}\right)^{2}\right)$
@@ 4.5 Uncertainty Approximation
We interpret these aggregated squared gradients in ‘hessian‘ as proxies for uncertainty. High squared partial derivatives indicate that small changes in the deformation field cause large changes to the rendered color [17, 24, 18], suggesting the reconstruction at that location is sensitive and potentially uncertain [11, 10]. Conversely, near-zero partial derivatives indicate that changes in local deformation parameters barely affect the color, suggesting more stable or confident estimates. Given the final accumulated $\mathbf{H}$ stored in our 3D grid, we define an uncertainty function $\sigma$ over the domain:
@@ 5.2 Evaluation Metrics
[…]
@@
    where $\hat{d}(u,v)$ is the predicted depth, $d^{(}u,v)$ is ground-truth depth, and $\Omega(p)$ is the set of remaining pixels after removing a fraction p of the domain based on uncertainty sorting. Plotting $\Delta\text{MAE}$ against the fraction of pixels removed quantifies how well our uncertainty measure correlates with actual reconstruction errors [14, 1]. If the model correctly identifies ”most uncertain” pixels as having high error, $\Delta\text{MAE}$ decreases rapidly when those pixels are removed.
@@
The figures show three colored lines for each dataset/view: the purple line represents sorting by BayesSDF’s uncertainty in descending order, the blue line represents sorting by variance or alternate uncertainty measure (e.g., an ensemble-based variance), and the red line represents the difference between the two curves’ $\Delta\text{MAE}$. When the red line reaches a low value, it implies that our Hessian-based uncertainty is identifying the same high-error pixels that the variance-based approach finds.
@@ 5.3 Analysis
These visualizations serve as a powerful diagnostic for identifying challenging areas and potential modeling failures. In the sparsification curves (Figure $\Delta\text{MAE}\text{ vs. }\text{Pixel Sparsification}$), both BayesSDF’s purple line and the variance-based blue line converge to similar cumulative MAE values as the sparsification approaches the least uncertain pixels. This convergence suggests that BayesSDF’s uncertainty ranking effectively captures error-prone pixels, similar to variance-based or ensemble-based uncertainty techniques [17, 24, 57, 18]. We present a table that summarizes the Area Under the Sparsification Error curve (with Mean Absolute Error) across multiple scenes:
@@
  • Despite not always having the absolute lowest metric, BayesSDF’s performance remains highly competitive while avoiding the large computational expense of ensemble training [11, 10].
@@
    Compact Integration into NeuS: BayesSDF only requires storing a Hessian-based statistic in the resolution-limited deformation grid, leading to a small parameter footprint while seamlessly integrating into the existing NeuS [47] rendering pipeline.
To validate that BayesSDF’s uncertainty maps correspond to poor geometry, we compare high-uncertainty regions against ground-truth geometry across all datasets [9, 35, 48]. We observe that high Hessian-based uncertainty consistently aligns with areas exhibiting large depth errors or missing surface detail such as reflective regions, occlusions, or sparse view coverage. These findings suggest that our Laplace approximation effectively identifies instability in the underlying SDF geometry [9, 35, 29, 54], making the uncertainty not just a confidence metric but a reliable indicator of physical reconstruction error.
@@ 6 Applications
The BayesSDF framework offers a transformative approach to uncertainty quantification, with significant implications for several domains that require precise and reliable 3D reconstructions. Below, we discuss the core applications of BayesSDF, emphasizing its technical strengths and relevance across various fields. BayesSDF is particularly well-suited for dense 3D scene reconstruction, a critical task in computer vision and robotics [20, 23, 26, 47]. BayesSDF quantifies spatial ambiguities in reconstructed regions. These uncertainty maps enable robotic systems to detect areas of high ambiguity, invaluable for autonomous navigation in dynamic or cluttered environments [53, 58, 41, 25]. For example, in Simultaneous Localization and Mapping (SLAM) systems, BayesSDF can improve loop closure detection by flagging uncertain regions in the map that require additional measurements [22, 50, 39]. Physics-based simulations in fields such as engineering and medical imaging rely on accurate 3D models [1, 14, 2] to simulate phenomena such as fluid dynamics or stress distribution [36, 42]. BayesSDF enhances these simulations by providing geometric models with quantified uncertainty [10, 11, 18]. For instance, in medical imaging, BayesSDF could quantify the variability in 3D reconstructions of anatomical structures, such as bones or blood vessels, to assess confidence in diagnostic or treatment planning tools [51, 17, 24]. BayesSDF opens new avenues in scientific research fields that depend on precise 3D modeling, such as computational biology and geology [33, 59, 60]. For example, in paleontology, BayesSDF can assist in reconstructing fossilized remains with quantified confidence levels [13, 29, 54], allowing researchers to distinguish between highly reliable and uncertain reconstructions.
@@ 7 Conclusion
We introduce BayesSDF, a probabilistic framework designed for uncertainty quantification in SDF-based 3D reconstruction models [29, 26, 47]. Motivated by scientific and physical simulation applications where accurate surface geometry is paramount [41, 19], BayesSDF integrates a Laplace approximation technique [35, 9, 24] over signed distance fields to quantify local reconstruction instability. Prior work has established that SDFs yield smoother, more consistent surfaces than NeRFs [20, 54, 28], and our method builds on this advantage to provide meaningful geometric uncertainty [14, 2, 1]. The result is a compact, efficient uncertainty estimator that aligns uncertainty directly with surface quality [18, 11]. By utilizing models like NeuS’ volumetric SDF space [47, 16], BayesSDF provides uncertainty maps that are inherently aligned with the reconstructed surfaces, enabling reliable and precise uncertainty quantification. The technical contributions of BayesSDF include a deformation field model based on hierarchical hashing [23], a computationally efficient Hessian-based curvature estimation [43], and ensemble-based epistemic uncertainty metrics [17, 57]. Together, these components address critical challenges in 3D uncertainty estimation, including the misalignment of uncertainty with surface geometry [27, 56] and the high computational costs associated with traditional sampling-based approaches [45, 8]. Through rigorous experiments on synthetic and real-world datasets [4, 33, 49], BayesSDF demonstrates outstanding performance, achieving very high calibration results [39, 51] and reduced computational overhead. Looking forward, the principles of BayesSDF can be extended to broader applications, such as dynamic scene understanding [30, 32, 25, 46] and hybrid models that combine SDFs with density-based neural representations [yariv2021volume, 7]. Moreover, the integration of BayesSDF with hardware-accelerated systems and distributed processing frameworks [55, 15] presents exciting opportunities for real-time 3D uncertainty quantification.